If cost function is C=wL + rK. Specify quantities of capital and labour that minimise costs?
1
Expert's answer
2021-07-26T17:55:02-0400

Solution:

C = wL + rK

The cost minimizing quantity of Labor (L):


"L = \frac{C - rK}{w}"

 

The cost minimizing quantity of Capital (K):


"K = \frac{C - Lw}{r}"